what is the equation of a circle tangent to the y-axis with center (-7,8) in standard form?
Answer by venugopalramana(3286)

what is the equation of a circle tangent to the y-axis with center (-7,8) in standard form?
Y AXS IS TANGENT TO CIRCLE.
THAT IS DISTANCE OF CENTRE OT THE CIRCLE TO Y AXIS=RADIUS OF THE CIRCLE.
BUT
DISTANCE OF CENTRE OT THE CIRCLE TO Y AXIS =X COORDINATE OF THE CENTRE=-7 AS CENTRE IS GIVEN AS (-7,8)
HENCE RADIUS =-7
STANDARD FORM OF EQN.OF CRCLE IS
(X-H)^2+(Y-K)^2=R^2
WHERE (H,K) IS CENTRE OF CIRCLE,AND R IS RADIUS
SO EQN. OF THE CIRCLE IS
(X+7)^2+(Y-8)^2=(-7)^2=49